OpenStudy (steve816):

Anyone know how to do this circular motion problem? I just wanna know how to start this.

OpenStudy (steve816):

A car is at rest at the top of a 80-meter high hill. It then rolls (accelerates) down a road that is inclined at 20° with respect to horizontal ground to the valley below. Once at the bottom of the hill, the driver of the car notices a curve in the road which causes cars to make a left turn. Cautiously, the driver applies the brakes for 2 seconds, resulting in the road exerting a friction force (μ = 1.0) on the car. After this, the car makes it through the turn perfectly (no friction needed for turning). The radius of the curve is 100 meters. At whant agle is the road banked (with respect to horizontal ground) throughout the turn?

OpenStudy (steve816):

Ok, I wanna know how to get the velocity upon reaching the curve...

OpenStudy (mrnood):

OK this is complex problem with a number of concepts - lets work through it First - dont worry about th emass of th ecar - it wiull cancel out in the end First phase - rolling down the slope. PE = mgh KE = mv^2/2 So the angle of th eslope seems irrelevant OK so far?

OpenStudy (mrnood):

Second phase - braking F - mu N N= mg Impulse = Ft = change of momentum = m (v2-v1) So oyu have v on entering the curve

OpenStudy (mrnood):

Third phase - circular motion Centripetal force = mv^2/r Centripetal force = N sin theta = mg sin theta I think that gets you there....

OpenStudy (mrnood):

TYPO Second phase should read F= mu N

OpenStudy (steve816):

Ok, I will try to work out the problem based on what you said. Thanks.

OpenStudy (steve816):

So for the first step, what equation would I use to find the velocity upon reaching the bottom of the hill?

OpenStudy (mrnood):

I gave it to you PE = mgh KE = mv^2/2 but don't try to solve numbers at each stage - you need to work through to the end to eliminate the mass

OpenStudy (steve816):

Ok, I don't quite understand how to relate PE to KE, is it possible to find velocity using kinematics?

OpenStudy (steve816):

Wait, nevermind, I get what you're saying.